# Partial Condition Numbers for Double Saddle Point Problems

### v1

@@ 6 Numerical experiments
The results in Tables 1-4 show that the upper bounds of the NCN, MCN and CCN provide very sharp estimates of the exact NCN, MCN and CCN, respectively. Additionally, it is observed that both the MCN and CCN, along with their upper bounds, are at most two orders of magnitude larger than the actual relative forward errors, offering more accurate estimates compared to the NCN and its upper bounds. These numerical results highlight the effectiveness of the proposed CNs and their corresponding upper bounds.
@@
$A=[(2ZZ^{T}+\Sigma_{1})\oplus\Sigma_{2}\oplus\Sigma_{3}]\in{\mathbb{R}}^{n\times n},\leavevmode\nobreak\ B=\left[\begin{matrix}N&-I_{2\tilde{q}}&I_{2\tilde{q}}\end{matrix}\right]\in{\mathbb{R}}^{m\times n},\leavevmode\nobreak$
@@ 7 Conclusion
Both authors contributed equally to the conceptualization, writing, and review of this work.

### v2

@@ 6 Numerical experiments
Additionally, it is observed that both the MCN and CCN, along with their upper bounds, are at most two orders of magnitude larger than the actual relative forward errors, offering more accurate estimates compared to the NCN and its upper bounds. These numerical results highlight the effectiveness of the proposed CNs and their corresponding upper bounds.
To evaluate the robustness of the proposed framework for computing the partial NCN, MCN and CCN of the DSPP with respect to small perturbations in the input data, we consider a perturbed coefficient matrix defined as $\widehat{\mathfrak{B}}=\mathfrak{B}+10^{-4}\Delta\mathfrak{B}$, where the block matrices of $\Delta\mathfrak{B}$ are specified as follows:
${\Delta}A=e_{p}\,{\tt std}(A)\,{\tt{randn}}(n,n),\leavevmode\nobreak\ {\Delta}B=e_{p}\,{\tt std}(B)\,{\tt{randn}}(m,n),\leavevmode\nobreak\ {\Delta}C=e_{p}\,{\tt std}(C)\,{\tt{randn}}(p,m),$
${\Delta}D=e_{p}\,{\tt std}(D)\,{\tt{randn}}(m,m),\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ {\Delta}E=e_{p}\,{\tt std}(E)\,{\tt{randn}}(p,p),$
where ${\tt std}(A)$ denotes the standard deviation of $A.$ For the perturbed coefficient matrix, we consider the perturbed DSPP $\widehat{\mathfrak{B}}\widehat{\bm{w}}={\bf b}.$
Table 6: Partial NCN, MCN and CCN for the original and perturbed DSPP by varying $e_{p}$ for $q=4.$
Original DSPP Perturbed DSPP
$e_{p}$ $\mathfrak{K}^{(2)}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $\mathfrak{K}^{\infty}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $\mathfrak{K}^{\infty}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $\mathfrak{K}^{(2)}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $\mathfrak{K}^{\infty}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $\mathfrak{K}^{\infty}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$
$0.10$ $022.0808E+$ 02 $008.3003E+$ 00 $011.2487E+$ 01 $022.0809E+$ 02 $008.3003E+$ 00 $011.2487E+$ 01
$0.15$ $026.2669E+$ 02 $011.4536E+$ 01 $011.6732E+$ 01 $026.2669E+$ 02 $011.4536E+$ 01 $011.6732E+$ 01
$0.20$ $031.0318E+$ 03 $012.0297E+$ 01 $022.8747E+$ 02 $031.0319E+$ 03 $012.0297E+$ 01 $022.8742E+$ 02
[…]
Table 7: Partial NCN, MCN and CCN for the original and perturbed DSPP by varying $e_{p}$ for $q=6.$
Original DSPP Perturbed DSPP
$e_{p}$ $\mathfrak{K}^{(2)}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $\mathfrak{K}^{\infty}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $\mathfrak{K}^{\infty}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $\mathfrak{K}^{(2)}_{\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $\mathfrak{K}^{\infty}_{mix,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$ $\mathfrak{K}^{\infty}_{com,\bm{\varphi}}(\mathbf{H},{\bf b};\mathrm{L})$
$0.10$ $034.2282E+$ 03 $013.3861E+$ 01 $031.8403E+$ 03 $034.2283E+$ 03 $013.3861E+$ 01 $031.8396E+$ 03
$0.15$ $035.9136E+$ 03 $014.8040E+$ 01 $021.5365E+$ 02 $035.9136E+$ 03 $014.8041E+$ 01 $021.5366E+$ 02
$0.20$ $031.1828E+$ 03 $011.1395E+$ 01 $012.3067E+$ 01 $031.1829E+$ 03 $011.1395E+$ 01 $012.3067E+$ 01
[…]
The computed partial NCN, MCN and CCN for both the original and perturbed DSPP are presented in Tables 6 and 7 for $e_{p}=0.10,0.15,0.20$. The perturbation introduced in the coefficient matrices is of the order $\mathcal{O}(10^{-3})$. From these tables, we observe that the partial NCN, MCN, and CCN remain nearly identical across all cases. This demonstrates that the proposed expressions for the partial CNs are robust and exhibit low sensitivity to small perturbations in the input data.
@@
$A=[(2ZZ^{{\bm{\top}}}+\Sigma_{1})\oplus\Sigma_{2}\oplus\Sigma_{3}]\in{\mathbb{R}}^{n\times n},\leavevmode\nobreak\ B=\left[\begin{matrix}N&-I_{2\tilde{q}}&I_{2\tilde{q}}\end{matrix}\right]\in{\mathbb{R}}^{m\times n},\leavevmode\nobreak$
@@ 7 Conclusion
Sk. Safique Ahmad: Contributed to the conceptualization and design of the study, provided critical feedback and suggestions on earlier versions of the manuscript, contributed to improvements, and reviewed and approved the final version of the manuscript. Pinki Khatun: Led the conceptualization and design of the study, drafted the initial version of the manuscript, developed and tested the software, provided comments on earlier versions, contributed to revisions and improvements, and reviewed and approved the final version of the manuscript.